OpenStudy (anonymous):

Solve the system of equations -2x − 9y = -6 x − y = 0

OpenStudy (anonymous):

-2x - 9y = -6 x - y = 0 are those the equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

with what method?

OpenStudy (anonymous):

I have found the solution to the problem but I need some help using the substitution with x+6y=8 5x+22y=31

OpenStudy (anonymous):

isolate x since it is the easiest x = 8 - 6y therefore, 5(8-6y) +22y =31 are you good to go?

OpenStudy (anonymous):

x=-9/8, no?

OpenStudy (anonymous):

no clue, too lazy to solve it

OpenStudy (anonymous):

The question requires me to put the values of x and y back into the equations to check to make sure the solutions are correct. I seem to be doing something wrong, because I am not getting the correct outcomes. Any trouble shooting tactics come to mind?

OpenStudy (anonymous):

subbing the values of x and y back in is the best way to check if it's correct. the only trouble shooting tactic would be to make sure not to make a calculation error. I got y = 6/11 try that